OpenStudy (anonymous):

The area of a circle is πr2. To find the area of a sector with a central angle of Θ, measured in radians, by what should you multiply πr2?

OpenStudy (anonymous):

A_circle = pi*r^2 A_sector = (1/2)*r^2*Θ do you see what it needs to be times by? if not I can help more

OpenStudy (anonymous):